maths89898:

Evaluate the expression 34 ÷ (14 − 5) × 2.

MidniteSunny:

By PEMDAS, the things in ze parenthesis goes first so 14-5

maths89898:

equals 9

maths89898:

then 9x2 right

MidniteSunny:

gud, next it would b exponents, but ur problem doesnt hav dat, and den multiplication and bingo chu got dat

MidniteSunny:

and yes, 9x2

maths89898:

equals 18

maths89898:

then 3x4=12

MidniteSunny:

where did chu get ur 3x4 ? ur problem says "34 ÷ " only

maths89898:

its an exponet

maths89898:

1 attachment
MidniteSunny:

o then dat changes everything

MidniteSunny:

first was parenthesis, 3^4 ÷ (14 − 5) × 2 14-5 = 9 next is exponents, 3^4 ÷ 9 × 2 3^4 =

MidniteSunny:

it's not 12 btw because for exponents, it's like 3^4 = (((3x3)3)3)

MidniteSunny:

@maths89898 make sense so far ?

maths89898:

@midnitesunny wrote:
it's not 12 btw because for exponents, it's like 3^4 = (((3x3)3)3)
Im confused on this part

MidniteSunny:

alright so u were taught how to solve exponents, yesh ?

maths89898:

Yes ive been taught but its very difficult

maths89898:

I thought it'd be 3+3+3+3

maths89898:

or 3x4

maths89898:

Okay so I searched it up and said it'd be 81

MidniteSunny:

nu, nu, it's more like (((3*3)3)3), or in other words the product of 3 times 3 multiplied by three, and then the product of that multiplied by three again

maths89898:

So 81 divided by 9 = 9

MidniteSunny:

we started the problem over because u didnt specify there was an exponent, so we did what was inside the parenthesis, then we solved the exponent, and since next on PEMDAS ish multiplication, we'll go to that: 81 ÷ 9 × 2 9 × 2 = ?

maths89898:

18!

MidniteSunny:

gud ! next is division, 81 ÷ 18 = ?

maths89898:

4.5

maths89898:

a decimal right

MidniteSunny:

mhm ^^

maths89898:

tysm for the help sunny!

maths89898:

^^

MidniteSunny:

ofc ! :D